The shortest side of a triangle is 8 inches less than the medium side, and the longest side is 4 inches longer than 3 times the shortest side. if the perimeter is 27 inches, find the length of the sides.
Answer by rfer(16322)

x-8+x+3(x-8)-4=27
x-8+x+3x-24-4=27
5x=63
x=12.6
x-8=4.6
3*4.6-4=9.8